# Electrically tunable plasmonic metasurface as a matrix of nanoantennas

### 12 Feb 2024 v1

@@ 3.1 Fabrication process
Witness samples were included during the deposition of hafnia and ITO for the purpose of evaluating the films independently. The hafnia layer had a density of 9.61268 g/cm and a thickness of 7.76 0.258 nm, confirmed using X-ray reflectometry (XRR). The electronic properties of the ITO layer were determined using an Ecopia HMS-3000 Hall effect measurement system in Van der Pauw configuration. The optical properties of the deposited ITO and hafnia layers were extracted from ellipsometry measurements using a spectroscopic ellipsometer (UVISEL Plus, Horiba Scientific). See Section 1 under supplementary material for details.
@@ 3.2 Electro-optic test set-up and characterization techniques
where has been measured and determined as described above. See section 2 under supplementary material for a an example calculation of the phase shift from the raw data.
@@ 3.5 Optical modelling
In our optical model, the refractive index of hafnia was taken from experimental ellipsometry data (see section S1), the relative permittivity of Au and Cr were taken from Johnson and Christy [29] and Rakić et al. [30], respectively, and the voltage-dependent relative permittivity distribution of the ITO - obtained from our electrostatic calculations in Subsection 4.2 - was used. We used mesh mapping along the top boundaries of the perturbed region in ITO and swept it throughout the 3D domain of this region in a step of 0.1 nm. A tetrahedral mesh was used for the remaining domains.
@@ 4.4 Electrical characteristics of the hafnia layer
Figure 5(A) illustrates the voltage-dependent specific capacitance (C-V characteristics) measured for a reflectarray. The measurements exhibit repeatability and reveal distinct accumulation patterns in the capacitors of each column, identified by pin number. The raw capacitance measurements were offset on average by pF relative to the calibrated probes, leading to an overestimation of the specific capacitance by . This offset was subtracted from the data presented in Figure 5(A). More details on this point are given in Section S3 of the supplementary document. The resulting C-V characteristics highlight variations in specific capacitance among the columns, suggesting discrepancies that should ideally converge into a single curve. These deviations are attributed to variances in the hafnia layer.
@@ 4.5 Epsilon near zero in ITO
The deposited ITO layer exhibits an epsilon-near-zero (ENZ) state in the near-infrared (NIR) wavelength range, as confirmed by the analysis on witness samples (refer to Section 1 in the supplementary material for details). The device’s performance is significantly influenced by the ENZ wavelength of the ITO layer. This perturbation in ITO permittivity near the nanoantenna, induced by carrier refraction, is facilitated by voltage gating, allowing for carrier density modulation through accumulation and depletion. The increased carrier density leads to a blue-shift in during accumulation impacts the reflectarray’s reflection properties.
@@
To estimate of the ITO layer in the reflectarray, we used the parameters obtained in Section 1 of the supplementary information file. Given that the carrier concentrations measured via the Hall effect align well with derived from the ellipsometry fittings, we calculated using an average , where is the Drude frequency for the bulk calculated using Eq. (LABEL:Sequ:ncal) based on the carrier concentration measured via the Hall effect, and represents the Drude frequency for the bulk obtained from the ellipsometry fitting. Using Eq. (13), the anticipated range of is from m to m. As the MOS structures enter the accumulation region, the carrier density in the perturbed ITO layer increases (Fig. 5(B)), resulting in a blue-shift of therein. If the breakdown field of hafnia is sufficiently high, may transition through the resonance wavelength of the reflectarray, leading to a significant phase shift in the reflected beam [7]. Due to the low BDF of the hafnia layer here, it was not possible to apply a sufficiently large bias voltage for this regime to be attained.
@@ 4.6 Speed of operation
The response speed of the reflectarray is inherently limited by the time required for the MOS capacitors to reach full charge under accumulation, then to discharge when the bias voltage is reduced. The measured specific capacitances given in Fig. 5(A), the corresponding areas (cf. Section 4 in the supplementary information file), and the minimum and maximum pin resistances determined to be k and k (cf. Section 3 in the supplementary information file), were considered in estimating the speed of operation. Based on these parameters, the time constant of the capacitors was calculated as , resulting in values ranging from 1.98 to 2.5 ns, from which the maximum frequency of operation is determined to be 63 MHz (for 2.5 ns).

### 23 Feb 2024 v2

@@ 3.1 Fabrication process
Witness samples were included during the deposition of hafnia and ITO for the purpose of evaluating the films independently. The hafnia layer had a density of 9.61268 g/cm and a thickness of 7.76 0.258 nm, confirmed using X-ray reflectometry (XRR). The electronic properties of the ITO layer were determined using an Ecopia HMS-3000 Hall effect measurement system in Van der Pauw configuration. The optical properties of the deposited ITO and hafnia layers were extracted from ellipsometry measurements using a spectroscopic ellipsometer (UVISEL Plus, Horiba Scientific). See Section LABEL:sup-section:ITOandHafniaCha under supplementary material for details.
@@ 3.2 Electro-optic test set-up and characterization techniques
where has been measured and determined as described above. See section SLABEL:sup-section:phase_measurement_example under supplementary material for a an example calculation of the phase shift from the raw data.
@@ 3.5 Optical modelling
In our optical model, the refractive index of hafnia was taken from experimental ellipsometry data (see section SLABEL:sup-section:ITOandHafniaCha), the relative permittivity of Au and Cr were taken from Johnson and Christy [29] and Rakić et al. [30], respectively, and the voltage-dependent relative permittivity distribution of the ITO - obtained from our electrostatic calculations in Subsection 4.2 - was used. We used mesh mapping along the top boundaries of the perturbed region in ITO and swept it throughout the 3D domain of this region in a step of 0.1 nm. A tetrahedral mesh was used for the remaining domains.
@@ 4.4 Electrical characteristics of the hafnia layer
Figure 5(A) illustrates the voltage-dependent specific capacitance (C-V characteristics) measured for a reflectarray. The measurements exhibit repeatability and reveal distinct accumulation patterns in the capacitors of each column, identified by pin number. The raw capacitance measurements were offset on average by pF relative to the calibrated probes, leading to an overestimation of the specific capacitance by . This offset was subtracted from the data presented in Figure 5(A). More details on this point are given in Section SLABEL:sup-section:rawdata of the supplementary document. The resulting C-V characteristics highlight variations in specific capacitance among the columns, suggesting discrepancies that should ideally converge into a single curve. These deviations are attributed to variances in the hafnia layer.
@@ 4.5 Epsilon near zero in ITO
The deposited ITO layer exhibits an epsilon-near-zero (ENZ) state in the near-infrared (NIR) wavelength range, as confirmed by the analysis on witness samples (refer to Section SLABEL:sup-section:ITOandHafniaCha in the supplementary material for details). The device’s performance is significantly influenced by the ENZ wavelength of the ITO layer. This perturbation in ITO permittivity near the nanoantenna, induced by carrier refraction, is facilitated by voltage gating, allowing for carrier density modulation through accumulation and depletion. The increased carrier density leads to a blue-shift in during accumulation impacts the reflectarray’s reflection properties.
@@
To estimate of the ITO layer in the reflectarray, we used the parameters obtained in Section LABEL:sup-section:ITOandHafniaCha of the supplementary information file. Given that the carrier concentrations measured via the Hall effect align well with derived from the ellipsometry fittings, we calculated using an average , where is the Drude frequency for the bulk calculated using Eq. (LABEL:Sequ:ncal) based on the carrier concentration measured via the Hall effect, and represents the Drude frequency for the bulk obtained from the ellipsometry fitting. Using Eq. (13), the anticipated range of is from m to m. As the MOS structures enter the accumulation region, the carrier density in the perturbed ITO layer increases (Fig. 5(B)), resulting in a blue-shift of therein. If the breakdown field of hafnia is sufficiently high, may transition through the resonance wavelength of the reflectarray, leading to a significant phase shift in the reflected beam [7]. Due to the low BDF of the hafnia layer here, it was not possible to apply a sufficiently large bias voltage for this regime to be attained.
@@ 4.6 Speed of operation
The response speed of the reflectarray is inherently limited by the time required for the MOS capacitors to reach full charge under accumulation, then to discharge when the bias voltage is reduced. The measured specific capacitances given in Fig. 5(A), the corresponding areas (cf. Section SLABEL:sup-section:pinAreas in the supplementary information file), and the minimum and maximum pin resistances determined to be k and k (cf. Section LABEL:sup-section:pinResistance in the supplementary information file), were considered in estimating the speed of operation. Based on these parameters, the time constant of the capacitors was calculated as , resulting in values ranging from 1.98 to 2.5 ns, from which the maximum frequency of operation is determined to be 63 MHz (for 2.5 ns).